OpenStudy (anonymous):

points (k+9,-4k-3) and (-2,1). y intercept of 9

OpenStudy (owlcoffee):

What do you want to find with this data?

OpenStudy (anonymous):

value of k and equation

OpenStudy (anonymous):

@owlcoffee

OpenStudy (owlcoffee):

Equation of what? line? circle?

OpenStudy (anonymous):

equation in y=mx+b

OpenStudy (owlcoffee):

That is called the equation of the line. You see, we have to first calculate the slope of this line, which is calculated by the points given, of course, a y-intercept implies that the x-coordinate of the point is zero, so let's call the two points: \[p(k+9, -4k-3)\] \[m(0,9)\] Where the point "m" is the y-intercept of the line, we want to calculate the slope using the formula \(m=\frac{ y_2-y_1 }{ x_2-x_1 }\) and jus plotting in the info: \[m _{m,p}=\frac{ (-4k-3)-8 }{ (k+9)-0 }\] yielding the result: \[m _{m,p}=\frac{ -4k-11 }{ k+9 }\] And pltting this to the form \((y-y_o)=m(x-x_o)\) where "m" is the slope we already calculated and we can replace yo and xo with any point we want: \[(y-8)=\frac{ -4k-11 }{ k+9 }(x-0)\] Further simplifying: \[y=\frac{ -4k-11 }{ k+9 }x+8\]